Figure 1. During last night the BNS range has been slowly fluctuating between 45Mpc and 55Mpc. The BNS range trend look almost identical to the trend of SR MAR TY and of the DARM optical gain measured by Hrec. Whenver SR TY is low the BNS range and optical gain are low. This morning around 7:05UTC I have put the SR TX offset to zero, as that will make SR TY higher to keep the same value of the DCP. This seem have been effective at increasing the DARM optical gain, and making it stable, and at reducing the SR TY fluctuations.
However, it has been that great at making the BNS range better. It dropped at 08:00 UTC. The B1p QD 50MHz out of loop signals, (QD2 V and QD1 H) change mean value and start to have more glitches about at the same time.